OpenStudy (anonymous):

Which of the following sentences correctly uses the verb ver? Answer Él ven la zanahoria. Él veo la zanahoria. Él ve la zanahoria. Él vemos la zanahoria. (i chose D is that correct?)

OpenStudy (cathyangs):

It would be C, because 've' is the third person singular conjugation.

OpenStudy (anonymous):

Ven = ellos (they see) veo = yo (I see) ve = Él (he sees) vemos = nos (we see)

OpenStudy (anonymous):

c is the right one

OpenStudy (anonymous):

c is the correct.